Saturday, 22 August

Special Activities

11:00 – 12:00  Auditorium Maximum H 105

Emmy Noether Lecture:

Cathleen Synge Morawetz, Courant Institute, New York University, New York, USA Variations on Conservation Laws for the Wave Equation Chair: Irene M. Gamba, Austin, Texas, USA

\underline{14:00 – 18:00 H 104}

Berlin as a Centre of Mathematical Activity A workshop organized by G. Israel (Rome, Italy) and E. Knobloch (Berlin, Germany)

14:00 – 14:15 G. Israel, University of Rome, Italy Introduction

14:15 – 14:45 David Rowe, Universität Mainz, Germany The Berlin-Göttingen Rivalry, 1870-1920

14:45 – 15:00 Discussion

15:00 – 15:30 Ivor Grattan-Guinness, Middlesex University, England Weierstrassian analysis and Cantorian set theory in Britain and the USA, 1890-1910

15:30 – 15:45 Discussion
15:45 – 16:15 Coffee break

16:15 – 16:45 U. Bottazzini, Centro Linceo Interdisciplinare “B. Segre” Accademia dei Lincei, Rome, Italy Weierstrass’ school of analysis and its influence on Italian mathematics

16:45 – 16:50 Discussion

16:50 – 17:20 S. Demidov, Russian Academy of Sciences, Moscow, Russia Russian mathematicians in Berlin in the second half of the XIXth until the beginning of the XXth century

17:20 – 17:25 Discussion

17:25 – 17:55 A. Dahan, CNRS Paris, France Images croisées des centres mathématiques de Paris et de Berlin, dans les années 1860/80

17:55 – 18:00 Discussion

ICM’98 – Program  1